Please help me solve this problem using substitution 2x – y = -16
x – y = -10

Answer by rothauserc(4718)

We are asked to solve
2x – y = -16
x – y = -10
solve second equation for x
x = y - 10
substitute for x in the first equation
2*(y-10) -y = -16
2y-20-y = -16
y = 4
x = 4 -10 = -6
therefore
x = -6 and y = 4